OpenStudy (yamyam70):

If I would solve for x 4^2x +1 = 8 ^ x + 5 then it would be , 2 ^ 2(2x + 1 ) = 2 ^ 3 ( x + 5) since we have same base, 4x + 2 = 3x +15 x = 13 Is this correct?

ganeshie8 (ganeshie8):

\(\large \color{Red}{\checkmark}\)

OpenStudy (yamyam70):

Thanks :) @ganeshie8 :))

ganeshie8 (ganeshie8):

np :) good job !!!

OpenStudy (unklerhaukus):

more brackets world make it clearer

OpenStudy (yamyam70):

sorry to confuse you :)

OpenStudy (unklerhaukus):

one way to check is to sub your soltion back into the origional equation x=13 → 4^(2x +1) = 8 ^ (x + 5) 4^(2*13 +1) = 8 ^ (13 + 5) 4^(27) = 8 ^ (18) 2^54 = 2^54

OpenStudy (yamyam70):

I see, thanks :)

OpenStudy (unklerhaukus):

good work